# Open-World Task and Motion Planning via Vision-Language Model Inferred Constraints

### v2

@@ IV-B Grounding Continuous Constraints with a VLM
To properly implement this action, we must ensure that the placement pose $p$ of detach action obeys the description $d$ of being “stably on the table and upright”. To this end, we prompt a VLM to generate code to implement a test on the pose $p$ directly that outputs a Boolean value (and can thus be used as part of VLMPose):
1def test_poses(p) -> bool:
2 ontop_table_bounds = modify_pose_bounds_to_be_ontop _of_object(’mug’, ’table’)
3 mug_on_table = position_within_bounds(mug.pose, ontop_table_bounds)
4 upright_orientation = abs(mug.pose.roll) < 0.1 and abs(mug.pose.pitch) < 0.1
5 return mug_on_table and upright_orientation
Given such a function, we can implement the VLMPose($d,o,p$) predicate by simply calling this function and passing in the pose pose $p$ at which the mug object is being placed. The description $d$ is passed into the VLM to generate this function. Given this implementation on VLMPose, the TAMP system will be constrained to solutions that respect this continuous constraint, in line with the intent of the task. Although we focus on Boolean functions as action constraints, our approach can be straightforwardly applied to nonnegative functions as action costs to, for example, minimize the distance from a placement to a table edge.
We additionally found that it is helpful to have the VLM output continuous constraints corresponding to the goal description $g$ itself, and then use these to output constraints on each of the discrete actions. We then feed its output from this step as part of the prompts for it to output constraints on every other action with description $d$ and a constraint fluent requiring a VLM implementation. Additional details related to prompting and constraint implementation are described in Appendix A-F, and a full walkthrough of generating constraints on an example task is shown in Appendix A-B2.
V Experiments and System Demonstration
Our experiments measure how well our approach is able to solve tasks with natural-language goals and requiring satisfying a wide range of different constraints. We compare our approach quantitatively against several ablations and baselines in simulation, and then implement and demonstrate its capabilities on real-world hardware.
@@ V Experiments and System Demonstration
CaP 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
CaP-sample 100% 20% 20% 0% 0% 0% 0% 0% 0% 0% 14%
No VLM 100% 100% 100% 0% 0% 0% 0% 0% 20% 0% 32%
No sample 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
No disc. 100% 100% 100% 0% 0% 0% 0% 0% 20% 0% 32%
No cont. 100% 100% 100% 100% 10% 60% 70% 0% 0% 20% 56%
No back. 100% 90% 100% 60% 80% 100% 40% 30% 0% 0% 60%
OWL-TAMP 100% 100% 100% 60% 100% 100% 100% 100% 70% 90% 92%
TABLE I: Success rates on all tasks. We report the success rate (i.e., whether the task was successfully accomplished, as measured by manually-written checks), across 10 random seeds. We run a one-tailed z-test with $\alpha=0.1$ and bold all top-scoring success rates. The final column displays the average success rate across all trials on all tasks.
Tasks. We introduce the tasks we tested our approach and baselines on, with additional details and discussion in Appendix A. These tasks are all implemented within the RAVENS-YCB Manipulation environment adapted from previous work [7, 55]. The environment features a 6 DoF UR5 robot arm with a Robotiq 2F-85 gripper in front of a tabletop strewn with objects from the YCB dataset [56], as visualized in Figure 3. The $10$ tasks are listed here in approximate order of difficulty, where more difficult tasks involve satisfying a larger number and complexity of constraints.
@@
    Berry1: “put the strawberry onto the light-grey region at the center of the table”. The optimal solution to this task requires a sequence of 2 skills.
@@
    Citrus: “pack the citrus fruit onto the plate”. The optimal solution to this task requires a sequence of 4 skills.
@@
    Berry2: Same as Berry1, but requires moving an obstacle out of the way of the light-grey region. The optimal solution to this task requires a sequence of 4 skills.
@@
    BerryCook: “Cook the strawberry by putting it in the pan, then finally simply place it in the bowl. The strawberry should only be in the bowl at the end!”. The optimal solution to this task requires a sequence of 4 skills.
@@
    FruitSort: “Put all the fruit to the left of the line bisecting the table”. The optimal solution to this task requires a sequence of 6 skills.
@@
    Coffee: “I want to pour some coffee into the cup; can you set up the cup on the table so I can do this properly?” (requires reorienting the cup so that it is placed ‘right-side-up’ on the table). The optimal solution to this task requires a sequence of2 skills.
@@
    Mug1: “Setup the mug so it’s upright, then put whatever object that fits inside of it”. The optimal solution to this task requires a sequence of 4 skills.
@@
    Mug2: “Place cutlery inside the mug and then place the mug itself on the table near the condiment” (the mug’s opening is obstructed by a large orange, which must be moved out of the way). The optimal solution to this task requires a sequence of 8 skills.
@@
    Mug3: Same as CutleryInMug2, except in the initial state the mug is not obstructed by an orange, but rather has a ball stuck inside it. This must be removed by ‘pouring’ it out of the cup itself). The optimal solution to this task requires a sequence of 8 skills.
@@
    SoupPour: “Serve the fruits on the white mat (make sure the peach is to the right of the apple” and pour soup into the red container”. The white mat is originally obstructed by the soup can as well as a spam can, which must be moved out of the way to successfully place the fruits. The optimal solution requires a sequence of 10 skills.
@@
    CaP: A version of Code as Policies [35] adapted to our domain, and using a VLM instead of an LLM as in the original work. This approach attempts to query a VLM to directly specify a function that outputs a full plan given an initial state. We then execute each action directly without any sampling. We provide the VLM with the same few-shot examples, as well as all the same helper functions that our approach uses.
@@
    No VLM: Inspired by Liu et al. [57] and Xie et al. [13], we simply prompt an LLM to translate the given natural language goal to a symbolic goal expression, and call a TAMP system to achieve this goal.
@@
    CaP-sample: A version of CaP inspired by the method of Curtis et al. [7] where the VLM generates a function that generates a variety of different plans (where each plan differs in the continuous parameters provided to the skill). We sample plans from this function and return the first full plan that does not violate any robot constraints.
@@
    No sample: A version of OWL-TAMP where we only allow one continuous sample per skill in the plan.
@@
    No cont.: A version of OWL-TAMP where we ablate the continuous constraint generation module (i.e., all continuous constraints come directly from the underlying TAMP system) but enable discrete constraint generation.
@@
    No disc.: A version of OWL-TAMP where we ablate the discrete constraint generation module but enable continuous constraint generation.
@@
    No backtrack: A version of OWL-TAMP where we ablate the underlying TAMP system’s ability to backtrack over different task skeletons (i.e., it is confined to attempting to refine the first plan skeleton it finds).
Experimental Setup. We use GPT-4o [1] as our VLM for all approaches. We report results over 10 independent random seeds for each task; we keep the natural language goal the same, and randomize the poses of a predetermined number of objects. For ‘CaP-sample’, we provide a sampling budget of $2500$ samples from its generated LMP for each task. For all other methods (except ‘CaP’ and ‘No sample’, where the sampling budget is set to $1$), we provide a sampling budget of $500$ per action in the plan. For methods that perform backtracking (i.e., all variants of our approach except ‘No back.’), we provide a backtracking budget of $5$ (i.e., we allow the method to attempt $5$ distinct plan skeletons). Our main metric of interest is each method’s success rate across tasks. We consider that a particular method has succeeded at a task if the robot does not violate any environmental constraints, and achieved the natural-language goal as judged by manually-written task-specific success detector functions. We also measure and report the ‘soundness rate’, which indicates how often an approach produces false positives (i.e., it believes it has accomplished a task, but actually has not). We report and analyze additional metrics on this domain, such as the total the wall-clock time, number of plan skeletons, and number of calls to a pretrained model to find a solution, in Appendix A-C. For all methods in the RAVENS-YCB Manipulation domain, we use a common TAMP system based on the search-then-sample TAMP system widely used in previous work [8, 12, 16, 17].
Results and Analysis. Table I shows our main success rate results for the RAVENS-YCB Manipulation domain. OWL-TAMP achieves the highest success rate of all methods in 9 out of the 10 presented tasks. The ‘No VLM’ baseline performs well in the simplest 3 tasks, where the natural language goal can be exactly translated into a goal expression in terms of predicates available to the underlying TAMP system, but fails almost completely in the remaining 7 tasks. Both ‘CaP’ and ‘No sample’ fail across all tasks. Upon inspection, we found that the discrete plan skeletons produced by these methods were generally correct, but that they were unable to select continuous parameters to satisfy the various continuous constraints within the domain. ‘CaP-sample’ performs better than ‘CaP’ due to its ability to sample different continuous values until it finds a satisfying set, but fails in tasks where the goal expresses non-trivial constraints because it does not explicitly attempt to generate and solve for these constraints. The ‘no cont.’ and ‘no disc.’ ablations of OWL-TAMP succeed in the first 3 tasks that do not require any discrete or continuous constraints beyond those embedded in the TAMP system, but struggle beyond these. Finally, the ‘no back.’ ablation of OWL-TAMP succeeds in most tasks except the final 4, which tend to require moving obstacles out of the way in order to satisfy constraints. We find that the VLM is able to generate appropriate discrete constraints to enable this in short-horizon tasks without much clutter (e.g. Berry2), but requires the assistance of a TAMP system equipped with backtracking to handle longer-horizon tasks featuring more objects.

### v4

@@ 3.1 Task and Motion Planning
The ${\texttt{Motion}}(q_{1},\tau,q_{2})$ constraint checks that $\tau$ is a valid trajectory connecting configurations $q_{1}$ and $q_{2}$, while ${\texttt{Kin}}(q,o,g,p)$ checks whether configuration $q$ satisfies a kinematic constraint for grasping object $o$ at placement pose $p$ with grasp $g$. The ${\texttt{AtGrasp}}(o,g)$ predicate indicates that object $o$ is currently grasped with grasp pose $g$. Additionally, for some of these constraints, we have standard robotics samplers that generate continuous values that satisfy these constraints, namely an inverse kinematics solver for Kin and motion planner for Motion. The challenge in planning is refinement: generating input parameter values for these actions (e.g. the pose p and grasp g for attach).
The goal in TAMP is typically specified by a symbolic goal $G$, a set of literals that must hold in the final state. The planner searches for a sequence of grounded actions that results in a state $s_{*}$ where all literals in $G$ evaluate to true. This search is conducted within a 3D model of the environment and objects: in real-world environments, this is constructed from the robot’s initial sensory observations such as point clouds. A key feature of TAMP is that the planner is capable of not just considering a single action sequence to achieve the goal (e.g., directly attach-ing then detach-ing the banana in Figure 1c), but rather exploring multiple sequences (via backtracking over an abstract plan) across multiple possible action sequences (e.g., first moving the milk carton out of the way) until it finds one that accomplishes the task.
4 Open-World Actions and Predicates
Traditional TAMP systems rely on a fixed set of symbolic predicates and operators, which limits their applicability to tasks that can be fully specified within this vocabulary. However, many open-ended tasks in unstructured environments involve object semantics, spatial relationships, or context-dependent goals that are not explicitly encoded in the domain definition. To address this, we propose a new abstraction that extends TAMP representations through open-world actions, which augment standard TAMP actions with natural language parameters, and open-world predicates, whose behavior is dynamically grounded via natural language.
Open-World Actions. An open-world action extends a parameterized TAMP action (Section 3.1) with an additional discrete argument: a natural language description $d$ as a string. This description specifies task-specific intent that constrains the set of valid continuous parameter values, such as object poses or robot configurations. Open-world actions enable novel behaviors that go beyond what can be expressed using the fixed symbolic vocabulary of a traditional TAMP system. For example, in Figure 1, the detach action is invoked with the description detach(“place the banana near the apple and pear”,  …), which narrows the feasible placement region from the entire table to a localized region, as visualized by the colored regions. A different description, such as “place the banana atop the orange” would restrict the final pose of the banana to be stably resting on the orange.
Open-World Predicates. To enforce that the intent expressed by the language description $d$ is achieved, we introduce open-world predicates, which are appended to the constraint list (con) of an open-world action. Traditional predicates rely on hand-engineered, fixed classifiers to evaluate whether a literal — a predicate with fully specified parameter values — is considered true or false. In contrast, an open-world predicate uses a classifier that is generated dynamically at planning time, conditioned on the description $d$ and other parameters (Figure 1b). We assume that these predicates take in $d$ as a parameter, along with any other objects and continuous parameters, and output a boolean value indicating that the parameters are set such that $d$ is achieved. We specifically use a VLM to generate this classifier function in the form of Python code (Section 5.2). The resulting classifiers act as constraints during planning, restricting the set of valid continuous parameters (e.g., object poses) according to whether they satisfy the task-specific intent described by $d$.
Example. We make the detach action open-world by modeling the description $d$ as a parameter and including the open-world predicate VLMConstraint in its constraints:
[…]
con: [Kin(q,o,g,p)(q,o,g,p), VLMConstraint(d,o,p)(d,o,p)]
pre: [AtGrasp(o,g)(o,g), AtConf(q)(q),
[…]
Let the open-world predicate ${\texttt{VLMConstraint}}(d,o,p)$ be true if placing object $o$ at placement pose $p$ satisfies description $d$. This allows the detach action to jointly respect both task-specific constraints specified by $d$ and engineered constraints such as kinematic feasibility (Kin). In Figure 1c, the TAMP planner rejection samples the set of reachable poses (generated by samplers for Kin) against VLMConstraint, implemented by the VLM-generated classifier function test_banana_pose, to yield a collision-free banana placement that is near the other fruit.
The detach action serves as a simple illustrative example that operates over a single object. Open-world actions can more generally use any number of objects and parameters, and model a variety of other interaction types. The following transport action models moving an object $o$ through pose waypoints $p_{1}$ and $p_{2}$ to then interact with object $o^{\prime}$. The waypoints $p_{1}$ and $p_{2}$ are determined by a VLM constraint VLMConstraint2 and can, for example, represent both the start and end of a pouring motion involving a cup (e.g. in the SoupPour task in Section 6) into a bowl $o^{\prime}$ as well as filling a cup $o$ from a faucet $o^{\prime}$. The key takeaway is that open-world actions enable planning with both manually engineered TAMP constraints and dynamically generated, task-specific constraints grounded in natural language, expressed as open-world predicates.
[…]
$p_{1}{:}\;\text{pose},\;p_{2}{:}\;\text{pose},\;o^{\prime}{:}\;\text{obj},q_{1}{:}\;\text{conf},\;q_{2}{:}\;\text{conf})$
con: $(q_{1},o,g,p_{1}), Kin$$(q_{2},o,g,p_{2})$,
[…]
5 Open-World Language-Based TAMP
Given a natural language instruction, images of the initial scene, a set of objects segmented within a scene, a TAMP system, and a collection of open-world actions and predicates, OWL-TAMP produces a valid plan that satisfies both manually engineered symbolic constraints and dynamically generated constraints from a VLM. Our approach operates in three stages: (1) it prompts a VLM to generate a discrete plan sketch of open-world actions relevant to the task (Section 5.1); (2) it grounds open-world predicates in the sketch into executable continuous constraints via VLM-generated classifier functions (Section 5.2); and (3) it solves the resulting TAMP problem using an off-the-shelf planner (Section 5.3).
5.1 Generating Open-World Action Ordering Constraints
Given the language instruction $L$ and initial images $\mathcal{I}$ (Section 3), we first prompt a VLM to produce a plan sketch $\hat{\pi}$: a partial sequence of open-world actions with discrete object arguments and natural language descriptions that reflect how to achieve the task specified by $L$. This sketch encodes a high-level plan structure, specifying a minimal set of actions and their order for a plan to be valid. In the banana task (Figure 1), the sketch could be:
$\hat{\pi}$ = [detach(“place the banana near the apple and pear”,  banana, …)].
In the weigh and dispose shortest object task (Figure 2c), the sketch might be:
$\hat{\pi}{=}$[detach(“weigh block on scale”, block, …),  detach(“place block in bin”, block, …)],
allowing us to detach the same object on different support surfaces. However, VLM-generated sketches are not guaranteed to be complete – they may omit necessary actions (e.g., forgetting to attach before detach) or propose actions that are infeasible due to low-level constraints such as collisions or kinematic feasibility. In the banana task (Figure 1), the VLM may omit a required attach(banana, …) before detach, or fail to consider that the milk carton obstructs all grasps on the banana.
To address these issues, we treat the plan sketch as a set of discrete action-ordering constraints. Specifically, we constrain the TAMP planner to generate plans that include $\hat{\pi}$ as a subsequence – that is, the specified actions must appear in the given order, though additional actions may be inserted before, between, or after them to ensure the plan is feasible.
We generate plan sketches using a few-shot prompting approach with a VLM with no fine-tuning, and describe how they are symbolically encoded into a TAMP system in Appendix A.1.2.
To enable a VLM to generate a plan sketch for a given task, we associate a natural language description of each available action with that particular action. Although we could directly prompt a VLM for relevant actions and goals, without a list of candidates, the VLM is likely to be syntactically and semantically inaccurate. Instead, we first ground the set of reachable actions $A$ and literals $\Lambda$ (i.e., grounded predicates) available to the TAMP system before prompting the VLM to return values in these sets. We use relaxed planning [6] from the initial state $s_{0}$ to simultaneously ground [24, 13] and explore the sets of reachable actions $A$ and literals $\Lambda$. We use placeholders to instantiate continuous parameters, such as optimistic values [19], to ensure a finite set of actions are instantiated. Similarly, we use placeholders for description parameters.
Algorithm 1 VLM Task Reasoning
1:procedure vlm-task-reasoning($s_{0},{\cal A},{\cal I},L$)
2:  $A\leftarrow\textsc{ground-actions}(s_{0},{\cal A})$
3:  $\Lambda\leftarrow s_{0}\cup\{l\mid a\in A.\;l\in a.{\it eff}\}$
[…]
5:  for $i\in[1,n-1]$ do
6:   $a_{i}.{\it eff}\leftarrow a_{i}.{\it eff}\cup\{{\texttt{Executed}}(i)\}$
7:   $a_{i+1}.{\it pre}\leftarrow a_{i+1}.{\it pre}\cup\{{\texttt{Executed}}(i)\}$
8:  $a_{n}.{\it eff}\leftarrow a_{n}.{\it eff}\cup\{{\texttt{Executed}}(n)\}$
9:  $G\leftarrow\{l_{1},..,l_{k}\}$
10:  return $\textsc{solve-tamp}(s_{0},A,G\cup\{{\texttt{Executed}}(n)\})$
Algorithm 1 presents the VLM partial planning pseudocode. It takes in a TAMP problem $\langle s_{0},{\cal A},\mathcal{I},L\rangle$, where $L$ is a text goal description and $\mathcal{I}$ is a set of initial image observations of the scene. It first grounds the set of actions $A$ reachable from $s_{0}$ using ground-actions. Then, it accumulates the set of reachable literals $\Lambda$ by taking the effects of all actions $A$. Then, it prompts query-vlm for a partial plan $a_{1},...,a_{n},l_{1},...,l_{k}]$ using actions $a_{i}\in A$ and goal literals $l_{j}\in\Lambda$ that achieve $L$. Here, the literals $l_{j}\in\Lambda$ are intended to be a translation of the goal into atoms, while the sequence of actions is a way of achieving that goal. Importantly, we have the VLM fill in the description parameter $d$ for each of these actions. We transform the original problem to so that solutions admit the partial plan as a subsequence. Specifically, we create a predicate Executed that models whether the $i$th action in the plan was executed and add Executed to the effects of action $a_{i}$ and the preconditions of action $a_{i+1}$. Finally, we make the planning goal $G=\{l_{i},...,l_{m}\}\subseteq L$ and $\textsc{Executed}(n)$.
@@ 5.1 Generating Open-World Action Ordering Constraints
CaP 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0
CaP-sample 100% $\pm$ 0.0 20% $\pm$ 12.6 20% $\pm$ 12.6 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 14% $\pm$ 3.4
Direct trans. 100% $\pm$ 0.0 100% $\pm$ 0.0 100% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 20% $\pm$ 12.6 0% $\pm$ 0.0 32% $\pm$ 4.7
No sample 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0
No disc. 100% $\pm$ 0.0 100% $\pm$ 0.0 100% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 0% $\pm$ 0.0 20% $\pm$ 12.6 0% $\pm$ 0.0 32% $\pm$ 4.7
No cont. 100% $\pm$ 0.0 100% $\pm$ 0.0 100% $\pm$ 0.0 100% $\pm$ 0.0 10% $\pm$ 9.5 60% $\pm$ 15.5 70% $\pm$ 14.5 0% $\pm$ 0.0 0% $\pm$ 0.0 20% $\pm$ 12.6 56% $\pm$ 5.0
No back. 100% $\pm$ 0.0 90% $\pm$ 9.5 100% $\pm$ 0.0 60% $\pm$ 15.5 80% $\pm$ 12.6 100% $\pm$ 0.0 40% $\pm$ 15.5 30% $\pm$ 14.5 0% $\pm$ 0.0 0% $\pm$ 0.0 60% $\pm$ 4.9
OWL-TAMP 100% $\pm$ 0.0 100% $\pm$ 0.0 100% $\pm$ 0.0 60% $\pm$ 15.5 100% $\pm$ 0.0 100% $\pm$ 0.0 100% $\pm$ 0.0 100% $\pm$ 0.0 70% $\pm$ 14.5 90% $\pm$ 9.5 92% $\pm$ 2.7
Table 1: Success rates on all tasks. Entries report the mean success rate over 10 random seeds and the binomial standard error in percentage points. We bold methods that are not significantly worse than the best method in that column under a one-tailed two-proportion $z$-test ($\alpha{=}0.05$). The final column pools successes across all 10 tasks (n=100n{=}100).
5.2 Grounding Open-World Predicates with a VLM
After obtaining a plan sketch from a VLM, we ground each open-world predicate into a concrete constraint that can be evaluated during planning. These predicates, which appear in the open-world actions within the plan sketch $\hat{\pi}$, are grounded by dynamically generating a classifier function for each predicate separately once the predicate is instantiated with a specific natural language description. The resulting classifier operates over continuous-valued arguments and returns a boolean indicating whether a given literal (i.e., a predicate with fully specified arguments), is considered true or false. We refer to these classifier functions as continuous constraints.
[…]
5.3 Planning with VLM-Generated Constraints
Given the discrete plan sketch and continuous constraints generated by the VLM, we can now solve the TAMP problem via an off-the-shelf planning algorithm, such as SeSaME [35] or PDDLStream [14] (see Appendix A.1.1). The planner searches for a valid sequence of discrete actions, constrained by the plan sketch, and solves for continuous parameter assignments satisfying both manually engineered and VLM-generated constraints. In the banana task (Figure 1), the planner first attempts to attach and detach the banana directly, but fails to find collision-free grasps due to the obstructing milk carton. It then backtracks and finds a feasible plan that first moves the milk carton before grasping and placing the banana at a pose that satisfies VLMConstraint and is reachable.
Figure 3: Simulated RAVENS-YCB Tasks. OWL-TAMP (a) reorients the mug and infers only the fork fits inside; (b) removes the ball from the mug before putting the fork inside, then moves it closer to the mustard; (c) moves the spam aside to fit fruits on the mat before pouring soup into the bowl.
6 Experiments and System Demonstration
Tasks. We evaluate on ten open-world manipulation tasks, spanning a variety of skills including obstacle removal, sorting, reorientation, insertion, and pouring. All tasks implemented in the RAVENS-YCB environment on a UR5 arm, adapted from prior work [10, 53] (see Appendix A.2).
@@ 6 Experiments and System Demonstration
    Berry1: “put the strawberry onto the light-grey region at the center of the table”.
@@
    Citrus: “pack the citrus fruit onto the plate”.
@@
    Berry2: Same as Berry1, but requires moving an obstacle out of the way of the light-grey region.
@@
    BerryCook: “Cook the strawberry by putting it in the pan, then finally simply place it in the bowl. The strawberry should only be in the bowl at the end!”.
@@
    FruitSort: “Put all the fruit to the left of the line bisecting the table”.
@@
    Coffee: “I want to pour some coffee into the cup; can you set up the cup on the table so I can do this properly?” (requires reorienting the cup so that it is placed ‘right-side-up’ on the table).
@@
    Mug1: “Setup the mug so it’s upright, then put whatever object that fits inside of it”.
@@
    Mug2: “Place cutlery inside the mug and then place the mug itself on the table near the condiment” (the mug’s opening is obstructed by a large orange, which must be moved out of the way).
@@
    Mug3: Same as CutleryInMug2, except that in the initial state, the mug is unobstructed by an orange but instead contains a ball, which must be removed by ’pouring’ it out.
@@
    SoupPour: “Serve the fruits on the white mat (make sure the peach is to the right of the apple” and pour soup into the red container”. The white mat is originally obstructed by the soup can as well as a spam can, which must be moved out of the way to successfully place the fruits.
@@
    CaP: Application of Code as Policies [36], using a VLM to generate plan code directly from initial states. Few-shot examples and helper functions match OWL-TAMP.
@@
    Direct trans.: Prompting an LLM to translate natural language goals to symbolic goals (see Section 3.1), then solving with TAMP. This baseline is inspired by [38, 50]
@@
    CaP-sample: CaP variant inspired by [10], where the VLM samples multiple continuous values for skills and selects the first set that satisfies robot constraints.
@@
    No sample: A version of OWL-TAMP allowing only one continuous sample per skill.
@@
    No cont.: Ablation of OWL-TAMP removing continuous constraint generation (continuous values come purely from TAMP). This is inspired by [51].
@@
    No disc.: Ablation of OWL-TAMP removing action-ordering constraints.
@@
    No backtrack: Ablation of OWL-TAMP where only the first plan skeleton is refined.
Experimental Setup. We use GPT-4o [1] as the VLM across all methods. Results are averaged over 10 seeds per task, with initial object poses randomized and language goals fixed. CaP-sample uses a 2500-sample budget per task; other methods (except CaP and No-sample, which use 1-sample) have a 500-sample budget per action. Backtracking methods (all OWL-TAMP variants except No-backtrack) may attempt to refine up to 5 plan skeletons. Success is measured by task completion without constraint violation, as judged by manually written detectors; soundness rate (false positive rate) is also reported. Additional metrics (wall-clock time, number of skeletons, and model queries) are in Appendix A.4. All methods use the same standard search-then-sample TAMP system [46, 35, 43, 34], which we detail in Appendix A.1.1.
6.1 Results and Analysis
Table 1 reports success rates. OWL-TAMP achieves the highest success rate on 9 of 10 tasks, outperforming all baselines. Direct trans. succeeds on the 3 simplest tasks, where language goals directly map to existing TAMP predicates, but fails on harder tasks requiring language-conditioned constraint generation. CaP and No sample find correct discrete plans but fail to select continuous parameters satisfying constraints. CaP-sample improves with sampling but struggles when satisfying constraints requires coordinated reasoning across discrete actions. Ablations disabling discrete (No disc.) or continuous (No cont.) constraint generation succeed only on easier tasks, while No backtrack struggles on longer-horizon problems requiring obstacle removal. By generating both discrete and continuous constraints, OWL-TAMP solves all tasks without any task-specific modifications.
Figure 4 reports soundness rates. Direct trans. and No cont. often declare success despite violating the true task intent (a false positive). OWL-TAMP produces almost no false positives — only a single case across all tasks — demonstrating precise grounding of language into constraints. CaP-sample has a high soundness only because it simply fails on most tasks (14% success rate) and declares it fails.
Failures for OWL-TAMP occurred in BerryCook, Mug3, and SoupPour: in BerryCook, the VLM generated an incorrect continuous constraint requiring the strawberry to be inside the bowl and pan simultaneously; in Mug3, the sample budget was exhausted despite correct generated constraints; and in SoupPour, the VLM generated an unsatisfiable plan sketch. By far the most common failure mode was that observed in BerryCook: the system generated infeasible continuous constraints that were impossible to satisfy. While we analyzed OWL-TAMP as a single-shot planner, when deployed in a replanning policy it can use observations from execution or human feedback to augment the language instruction and adapt its behavior after failures.
6.2 Real-World System Deployment
In order to demonstrate our system’s applicability to open-world tasks in the real-world, we deployed OWL-TAMP on a dual-arm robot (Kinova Gen3 arms with a pan-tilt head camera) to perform 19 natural language manipulation tasks, listed below (Figure 2, full list in Appendix A). Please see our supplement for video of the robot’s execution.
  • $\bullet$
    “Put the orange and apple on the plate.”
  • $\bullet$
    “Place the strawberry and lime each in the bin that matches their color.”,
  • $\bullet$
    “Stack the blocks into a tower by increasing hue.”
  • $\bullet$
  • $\bullet$
  • $\bullet$
    “Put the orange where the apple is initially”.
  • $\bullet$
  • $\bullet$
  • $\bullet$
    “Put the green block between the blue and red ones”
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    “Fry two eggs at the front of the pan”
  • $\bullet$
    “Fry the spam on the pan and serve it on the plate”
  • $\bullet$
    “Weigh the shortest object and put it in the bin”
  • $\bullet$
    “Put the banana near where the other fruit are initially”
  • $\bullet$
    “Place the red block so that it’s aligned with the other two blocks”
In each task, the robot received the camera image and natural language command, estimated the scene geometry using a strategy similar to that in [9], generated constraints and plans via PDDLStream [14], and executed plans open-loop on hardware. No changes were made to the system across tasks. OWL-TAMP successfully completed all tasks, despite substantial variation in object configurations and goals. Many tasks required adapting grasps (e.g., side grasps over top-down), managing collisions and reachability, or backtracking to move obstacles. These results highlight OWL-TAMP’s ability to flexibly and robustly integrate vision, language, and planning without any task-specific tuning.
While our system was able to solve versions of each of these tasks, we observed several failures on similar tasks and variants. The primary failure mode of OWL-TAMP was the same as in the simulated experiments: generating incorrect continuous constraints (e.g. the VLM generated a constraint for an object to be in collision with another object). A less common but still prevalent failure mode was generating correct continuous constraints that could not be satisfied with our finite sampling budget (e.g. a particular placement region is unreachable because the real-world robot has a rather limited workspace). The majority of failures of the overall system were caused by common errors made by third-party real-world perception and control modules such as false negatives when segmenting objects and minor collisions due to imperfect trajectory tracking.